A person who is about to retire has accumulated $100,000 in a savings account. Suppose that the person withdraws $8,195.23 from the savings account at the end of each year for 20 years, at which time the account is totally depleted. What is the interest rate, based upon annual compounding?


1
Expert's answer
2022-02-14T14:46:53-0500

"PV=A\times\frac{1-\frac{1}{(1+r)^n}}{r}"


PV=100 000

A=8195.23

n=20

"100 000=8 195.23\times\frac{1-\frac{1}{(1+r)^{20}}}{r}"

"12.20=\frac{1-\frac{1}{(1+r)^{20}}}{r}"


"12.20=\frac{1-\frac{1}{(1+r)^{20}}}{r}"

"12.20=\frac{1}{r^{21}}"


"r^{21}=0.081952"

r=0.887697